OpenStudy (anonymous):

Divide 3x^3+3x^2+2x-2 by x+3 using long division. a) 3x^2-6x+20 b) 3x^2+12x-16 c)3x^2+12x-16, R 58 d) 3x^2-6x+20, R -62

OpenStudy (anonymous):

connections?

OpenStudy (anonymous):

Yeahh

OpenStudy (anonymous):

which one is this? lol

OpenStudy (anonymous):

Like lesson of what do you mean? haha

OpenStudy (anonymous):

lol yes the lesson :P

OpenStudy (anonymous):

Oh Lesson 4: Dividing Polynomials. Unit 5

OpenStudy (anonymous):

1-11 is B,B,A,B,D,A,C,B,D,B,C

OpenStudy (anonymous):

Omg you are amazing. lol Thank you!

OpenStudy (anonymous):

12. x^3 – 3x^2 – 10x =x(x^2–3x–10) =x(x^2–5x+2x–10 =x{x(x–5)+2(x–5)} =x(x–5)(x+2)

OpenStudy (anonymous):

13. 27x^3+125=0 27x^3=–125 x^3=–125/27 x^3=(–5/3)^3 x=–5/3

OpenStudy (anonymous):

got the rest?

OpenStudy (anonymous):

Not 14. But I have 15. you have completely made my day!

OpenStudy (anonymous):

okay ill give you 14

OpenStudy (anonymous):

P(-3) = 81 - 9(-27) - 5(9) + 9 + 4+4

OpenStudy (anonymous):

opps lol wait!

OpenStudy (anonymous):

14. P(-3) = (-3)^4 - 9(-3)^3 - 5(-3)^2 - 3(-3)+4 P(-3) = 81 - 9(-27) - 5(9) + 9 + 4+4 P(-3) = 81 + 243 - 45 + 9 + 4 P(-3) = 324 - 45 + 13 P(-3) = 337 - 453 P(-3) = 292

OpenStudy (anonymous):

Thank you soo so much

OpenStudy (anonymous):

welcome!